OpenStudy (anonymous):

y=2x+6 y=3x Solve the eqations using substution

OpenStudy (anonymous):

3x=2x+6 x=6

OpenStudy (anonymous):

x=6, y = 18

OpenStudy (anonymous):

3x = 2x+ 6... subtract 2x from both sides... x = 6... plug 6 into the y=3x equation yeilding y = 3(6)... y = 18

OpenStudy (lilg132):

y = 2x + 6 y = 3x 2x + 6 = 3x 6 = 3x - 2x 6 = x